Plot the points with polar coordinates (-5, 1) and 3, -* 3 using the pencil. / WE xs ? a 24
Answer #1

(-5, \pi) \mbox{ is a point on the x axis with radius 5}

(3, \frac{-5\pi}{3}) \mbox{ is a point on the line } \theta=\frac{\pi}{3} \mbox{ with radius 3}

• (-5, A) Red dot is (-5,pi) 01(3, -35) Blue dot is (3.-5pi/3)